Four resistors are connected in a circuit with a b
Answer #1

As we see, 30 Ω & 60 Ω resistors are in parallel they have same voltage across them !!

i1 * 30 =  i2 * 60
120 * 30 = i2 * 60
i2 = 60 mA

Current in 60 Ω resistor, i2 = 60 mA

Current in the 80 Ω Resistor = Current in 60 Ω resistor + Current in 30 Ω resistor
Current in the 80 Ω Resistor = (120 + 60 ) mA
Current in the 80 Ω Resistor = 180 mA

emf of the battery, = 60*(60/1000) + 80 * (180/1000)
emf of the battery, = 18 v

Current in 20 Ω Resistor, = 18/20 = 0.9 A

Power dissipated in 20 Ω Resistor, = 0.9^2 * 20 = 16.2 W

Current in the battery, = 0.9 + 180/1000 = 1.08 A